OpenStudy (anonymous):

if f(x)=x+3 what is f(g(x)

OpenStudy (anonymous):

where is g(x)? \(\large f(x)=x+3 \Longrightarrow f(g(x))=g(x)+3\) whatever g(x) is substitute into there

OpenStudy (anonymous):

thanks a lot

TheSmartOne (thesmartone):

\(\Huge\cal\color{blue}{Welcome~to~Open~Study!!}\) It looks like you are new here. This tutorial can help you get started. http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

TheSmartOne (thesmartone):

DId you get your question answered?

TheSmartOne (thesmartone):

I'll take that as a yes :)